# Integrating Low-Altitude SAR Imaging into UAV Data Backhaul

### v1

@@ VI Conclusion
In this paper, we presented a low-altitude SAR imaging framework that reuses native 5G NR OFDM waveforms and extends conventional pilot-based SAR to fully collect data symbols for sensing. To cope with the performance loss induced by discrete QAM data symbols, we embedded several tailored TF-domain filtering schemes into the conventional RD processing chain, which suppress the signaling randomness and render the echoes compatible with standard RD-based imaging. We further introduced the NMSE of a reference point-target’s profile as a more comprehensive metric that jointly reflects $\text{SNR}_{\text{out}}$, ISLR, and PEL. Simulation results confirm that the proposed data-aided imaging approach substantially outperforms the pilot-only baseline, highlighting its strong potential for integrating low-altitude SAR imaging into UAV data backhaul.
@@ Appendix A Proof of Lemma 1
We may exploit the correlation function of a given signal to characterize how fast it varies by introducing the notion of correlation time. Specifically, the correlation function of $r_{m}\left(k-k_{q,m}\right)$ is derived in (60), which is placed at the top of this page.
@@
As implied by $k_{q,m}=(\Delta R_{q,m}+\bar{R}_{q})/\rho_{r}$, the correlation time of envelope $r_{m}\left(k-k_{q,m}\right)$ can thus be inferred by letting
$k_{q,m+\Delta m}-k_{q,m}=1,$ (61)
or equivalently
$\Delta R_{q,\Delta m}=\rho_{r}.$ (62)
This leads to the approximated correlation time of envelope as
@@
For a typical low-altitude SAR imaging scenario operating in the sub-6 GHz band, the synthetic aperture length is on the order of a hundred meters, whereas the detection range can reach the kilometer scale. As a consequence, it is readily verified that
@@ Appendix B Proof of Theorem 1
where $\sum_{m}e^{j\frac{1}{2}\Phi^{\prime\prime}(\tilde{m})(m-\tilde{m})^{2}}\approx\sqrt{\frac{2\pi}{\left|\Phi^{\prime\prime}(\tilde{m})\right|}}e^{j\mathrm{sgn}\left(\Phi^{\prime\prime}(\tilde{m})\right)\frac{\pi}{4}}$ [cumming2005digital], which completes the proof.

### v2

@@ VI Conclusion
In this paper, we presented a low-altitude SAR imaging framework that reuses native 5G NR OFDM waveforms and extends conventional pilot-based SAR to fully collect data symbols for sensing. To cope with the performance loss induced by discrete QAM data symbols, we embedded several tailored TF-domain filtering schemes into the conventional RD processing chain, which reveals the mapping mechanism between the signaling randomness and the imaging quality. We further introduced the NMSE of a reference point-target’s profile as a complementary metric that jointly reflects $\text{SNR}_{\text{out}}$, ISLR, and PEL. Simulation results confirm that the proposed data-aided imaging approach substantially outperforms the pilot-only baseline, highlighting its strong potential for integrating low-altitude SAR imaging into UAV data backhaul.
@@ Appendix A Proof of Lemma 1
We may exploit the correlation function of a given signal to characterize how fast it varies by introducing the notion of correlation time. Specifically, the correlation function of $r_{m}\left(k-k_{q,m}\right)$ is derived in (60).
@@
As implied by $k_{q,m}=(\Delta R_{q,m}+\bar{R}_{q})/\rho_{r}$, the correlation time of envelope $r_{m}\left(k-k_{q,m}\right)$ can thus be inferred by letting $k_{q,m+\Delta m}-k_{q,m}=1$, or equivalently $\Delta R_{q,\Delta m}=\rho_{r}$. This yields the approximated correlation time of envelope as
@@
For a typical low-altitude SAR scenario at the sub-6 GHz band, the synthetic aperture length is on the order of a hundred meters, whereas the detection range can reach the kilometer scale. Consequently, it is readily verified that
@@ Appendix B Proof of Theorem 1
where $\sum_{m}e^{j\frac{1}{2}\Phi^{\prime\prime}(\tilde{m})(m-\tilde{m})^{2}}\approx\sqrt{\frac{2\pi}{\left|\Phi^{\prime\prime}(\tilde{m})\right|}}e^{j\mathrm{sgn}\left(\Phi^{\prime\prime}(\tilde{m})\right)\frac{\pi}{4}}$ [5], which completes the proof.